Please show work
Find the distance between the points (2, -2) and (-4, 3)

Thank you!

Found 2 solutions by venugopalramana, retaw04:
Answer by venugopalramana(3286)

Find the distance between the points (2, -2) and (-4, 3)
D=SQRT{(X2-X1)^2+(Y2-Y1)^2}={(2+4)^2+(-2-3)^2}^0.5=(36+25)^0.5=61^0.5

Answer by retaw04(1)

for easier explanation... distance formula is the sq.rt of (x2-x1)^2 + (y2-y1)^2
(-4-2)^2 + (3--2)^2 = (-6)^2 + (5)^2 = 36+25 = 61 = sq.rt of 61 ans.